# Universality of Efimov states in highly mass-imbalanced cold-atom mixtures with van der Waals and dipole interactions

### v1

@@ II.2 Quantum Defect Theory for non-dipolar system
The QDT parameter $K^{c}$ is related to the short-range phase, hence the short-range boundary condition. Indeed, at short distance, one can substitute the low-energy form of the wavefunction Eq. (28) because the energy is much smaller than the potential, and imposing $u_{\ell}(r=R_{\mathrm{min}})=0$, one finds
@@
the only difference with Eq. (2) is the presence/absence of the $1/r^{2}$ attraction. At short distance $r\sim R_{\mathrm{min}}\ll r_{\mathrm{vdw}}$, this attraction is much smaller than the van der Waals force so that the two equations are essentially the same. More physically, the short-range phase of the three-body system is dominated by the potential between the heavy atoms. This assumption should be particularly true for a highly mass-imbalanced system where the atomic polarizability of the heavy atoms, hence the van der Waals force and higher-order forces are much stronger than those for the light atom. Therefore, the QDT parameter of the three-body system can be approximated by that of two heavy atoms $K^{c}_{\mathrm{3body}}\simeq K^{c}_{\mathrm{2body}}$, which amounts to adopting the same short-range cutoff $R_{\mathrm{min}}$ value for the two- and three-body systems.
@@
Figure 1: Binding energy of the Efimov states with the unitary heavy-light interaction for (a) bosonic and (b) fermionic heavy atoms interacting via the van der Waals interaction. The mass ratio are (a) $M/m=27.5855...$ and (b) $M/m=27.7520...$, corresponding to 166Er-6Li and 167Er-6Li, respectively. The numerical results for the bosons (blue circles) and fermions (orange squares) with $\ell_{\mathrm{max}}=20$ are shown, together with the analytical curves of Eq. (7) with Eq. (8) (black solid) and Eq. (9) (green dashed). The vertical thin dotted lines (black) denote the thresholds at which the higher angular momentum bound states of $\ell=2$ ($\ell=3$) start to appear for the bosons (fermions), obtained from Eqs. (35) and (8).
@@ III.1 Non-dipolar case $a_{dd}=0$
Figure 2: Binding energy of the Efimov states plotted against (a)(b) the $s$-wave scattering length $a^{(\mathrm{HH})}$ for the bosons, and (c)(d) $p$-wave scattering volume $v^{(\mathrm{HH})}$ for the fermions, obtained by converting $R_{\mathrm{min}}$ in Fig. 1 into $a^{(\mathrm{HH})}$ and $v^{(\mathrm{HH})}$ using Eqs. (8) (13) (14). The left (right) pannels are negative (positive) scattering length/volume sides, which should smoothly connect with each other in the unitary limit $01/a^{(\mathrm{HH})}=$ 0, $01/v_{p}^{(\mathrm{HH})}=$ 0 corresponding to the border between the left and right panels. The solid curves (black) are the analytical results of Eqs. (15) (16).
First, we show the results without the dipole interaction: in Fig. 1, we show the binding energy of the Efimov states as the short-range boundary $R_{\mathrm{min}}$ is varied. As $R_{\mathrm{min}}$ gets smaller, the short-range van der Waals attraction increases, leading to the successive appearance of bound states and increase of the binding energy. The numerical data points almost perfectly agree with the analytical curves obtained with the QDT in Eq. (7), particularly at small energy. There are a few points which significantly deviate from the analytical curves. They are ascribed to higher angular momentum states of $\ell=2,4,...$ for the bosons and $\ell=3,5,...$ for the fermions, respectively. The thresholds at which the $\ell=2$ ($\ell=3$) bound states start to appear, denoted by the vertical dotted lines in Fig. 1 (a) (and (b)), agree with most of the stray data points, supporting this interpretation. A few stray points which agree with neither are ascribed to higher angular momentum channels $\ell\geq 4$ ($\ell\geq 5$). This classification of the states has also been confirmed from the shape of the wavefunctions. Because the higher angular momentum states are localized to shorter distances than the $s$-wave ($p$-wave) state owing to the centrifugal barrier, they are more sensitive to $R_{\mathrm{min}}$, resulting in the steeper binding energy curves. The binding energy curves of different partial waves do not render an avoided crossing but simply cross because the angular momentum is a good quantum number in the absence of the dipole interaction.
@@ III.3 Moderate dipole interaction $a_{dd}\sim r_{\mathrm{vdw}}$
In Fig. 5(g)-(l), we show the binding energies of the Efimov states plotted against $a^{(\mathrm{HH})}_{\mathrm{with-dd}}$. Similar to Fig. 5(a)-(f), the data with different values of $R_{\mathrm{min}}$ collapse into a single curve, demonstrating the universality. One notable difference is that the agreement between the numerical data and the purely van der Waals analytical curves (black) is much better in (g)-(l) than those in (a)-(f): except for the broad avoided crossing at $a_{\mathrm{with-dd}}^{(\mathrm{HH})}/r_{\mathrm{vdw}}\simeq 1$, a significant fraction of the numerical results agree well with the van der Waals curves not only for the weak dipole interaction, but also for the realistic dipole strength (Fig. 5(i)(j)) and stronger dipole interaction ((k)(l)). This suggests that the van der Waals quantum QDT of the bosonic Efimov states in Eq. (15) can capture the universal behavior of the Efimov states in the presence of the dipole interaction, once the effects of the dipole interaction are incorporated into the $s$-wave scattering length between the heavy atoms: $a^{(\mathrm{HH})}=a^{(\mathrm{HH})}_{\mathrm{with-dd}}$. In other words, the major role of the dipole interaction is to change, hence renormalize the value of the $s$-wave scattering length between the heavy atoms, so that the van der Waals universality of the Efimov states in Eq. (15) persists with the renormalized $s$-wave scattering length $a^{(\mathrm{HH})}_{\mathrm{with-dd}}$. This renormalized van der Waals universality for the dipolar bosonic Efimov states we have found here seems non-pertubative; the dipole interaction is so large that it significantly changes both the binding energies (Fig. 5(c)-(f)) and the $s$-wave scattering length (Fig. 6(b)(c))) from the purely van der Waals system. Some horizontal data points in Fig. 5(g)(l) which are not present in (a)-(f) are artefacts of the higher-partial-wave induced resonances: the two-body problem between the heavy atoms in Eq. (17) not only shows the $s$-wave dominated resonances, but also shows the resonant enhancement of $a^{(\mathrm{HH})}$ induced by higher-partial-wave channels [59, 61], whose approximate positions are indicated by the vertical black dotted lines in Fig. 6. The narrow nature of these higher-partial-wave induced resonances leads to an abrupt change of $a^{(\mathrm{HH})}_{\mathrm{with-dd}}$ as $R_{\mathrm{min}}$ is changed, resulting in flat energy spectra in Fig. 5(g)-(l). Equation (15) cannot explain these flat data points. We note however that these flat energy spectra are absent in Fig. 5(a)-(b), and they should therefore be interpreted as artefacts originaing from the two-body effect of $a^{(\mathrm{HH})}$ rather than a genuine three-body phenomenon. Physically, when we attempt to evaluate the binding energies of the Efimov states from an experimentally obtained value of $a^{(\mathrm{HH})}$ (see Sec. IV), we can simply use the van der Waals curves of Eq. (15) by neglecting the stray flat spectra, except for the case where the two Er atoms are coincidentally in the proximity of the $s$-wave resonance induced by a $d$-wave or higher-angular-momentum molecular state.
@@
In Fig. 8, we show the binding energies of the fermionic Efimov states plotted against $K_{c,F}^{(3D)}$ as obtained above. The numerical data of different short-range $R_{\mathrm{min}}$ values collapse into a single curve, demonstrating the universality of the fermionic Efimov states. The universality holds particularly well for a realistic $a_{dd}$ value of 167Er-167Er (Fig. 8(a)(b)), where only a few points of higher angular momentum character deviate from the universal curve. In contrast to the bosonic case, we cannot find an analytical expression which could explain the universal energy spectra owing to the impossibility of using $v_{p}^{(\mathrm{HH})}$ hence Eq. (16). The avoided crossing at $K_{c,F}^{(3D)}\simeq-1$ originates from the $\ell=3$ channel which is present without the dipole interaction at around $r_{\mathrm{vdw}}/v_{p}^{(\mathrm{HH})}\simeq-0.3$ (Fig. 2(c)). It shifts toward a larger negative $K_{c,F}^{(3D)}$ side with a decreased slope as the dipole strength is increased from a realistic value of 167Er-167Er (Fig. 8(a)(b)) toward larger values (Fig. 8(c)-(f)). The foothill of the avoided crossing shows a larger variance for stronger dipole interactions (Fig. 8(c)-(f)). This is a precursor of the breakdown of our assumption that the single-channel short-range phase in $\ell=1$ universally characterizes the three-body system (see Sec. III.4), especially for a very strong dipole interaction and large $K_{c,F}^{(3D)}$ (Fig. 8(c)(e)(f)).
@@ III.4 Strong dipole interaction $a_{dd}\gg r_{\mathrm{vdw}}$
The data points with a small slope in Fig. 9, including the weakly bound states, collapse into a single curve for both the bosonic and fermionic systems, suggesting that $K_{c}^{(1D)}$ universally characterizes the short-range phase, hence the Efimov states with the strong dipoles. On the other hand, the data points with steeper slopes do not seem to show the universality with respect to $K_{c}^{(1D)}$. Since those states have significant contributions from large angular momentum states $\ell\gg 1$ and the angular insensitive nature of their short-range phase can break down, it is natural that their features cannot be universally captured by the one-dimensional scattering model in Eqs. (20) (21) where the angular momentum is absent.
@@ IV Three-body parameters of Er-Er-Li Efimov states
166Er-6Li 75.5 65.5 68 0.495 -2.66 0.0712
0.107 -12.4 0.331
2.29×$\times 10^{-2}$57.3 1.54
4.94×$\times 10^{-3}$266 7.13
168Er-6Li 75.8 66.3 137 0.352 -3.74 0.100
7.65×$\times 10^{-2}$17.2 0.461
1.66×$\times 10^{-2}$79.0 2.12
3.62×$\times 10^{-3}$363 9.74
170Er-6Li 76.0 67 221 0.298 -4.40 0.118
6.55×$\times 10^{-2}$20.0 0.538
1.44×$\times 10^{-2}$91.3 2.45
3.16×$\times 10^{-3}$416 11.2
While there are spectroscopic methods to directly associate the Efimov trimers and observe their binding energies [74, 75, 76, 77], it is challenging to perform such experiments, especially in the unitary limit. It is more straightforward to observe the three-body loss rate of the atoms for variable $s$-wave scattering lengths $a^{(\mathrm{HL})}$, and search for the negative scattering length $a=a_{-}^{(\mathrm{HL})}$ at which the loss rate peak appears [11, 12, 13, 14, 16, 19]. Alternatively, one can prepare a mixture of heavy-light Feshbach molecules and heavy atoms, and observe the dimer-atom loss rate to find the positive scattering length $a=a_{*}^{(\mathrm{HL})}$ at which its peak appears [17]. The values $a_{-}^{(\mathrm{HL})}$ and $a_{*}^{(\mathrm{HL})}$ correspond to the points at which the Efimov states dissociate into three atoms, or a dimer and an atom, respectively. Notably, any of $a_{-}^{(\mathrm{HL})}$, $a_{*}^{(\mathrm{HL})}$, or $\kappa_{*}$ can be regarded as the universal scale characterizing the Efimov states because they can be related with each other by the universal Efimov theory [1, 2, 3, 4, 5, 6]. Using the zero-range Efimov theory, we have estimated $a_{-}^{(\mathrm{HL})}$, $a_{*}^{(\mathrm{HL})}$ from the $\kappa_{*}$ values. The results are presented in the last two columns of Table 1. The values of $a_{-}^{(\mathrm{HL})}$ corresponding to the most tightly bound Efimov states are likely to be either around $a_{-}^{(\mathrm{HL})}/r_{\mathrm{vdw}}\approx-3$-$4$ or $-12$-$20$. With the low-energy formula Eq. (15) and the zero-range theory, we cannot conclude which of them are the ground Efimov state. We also note that they can be significantly affected by finite-range effects owing to their relatively large binding energy, while results for the more weakly bound states in Table 1 should be more quantitatively reliable. Similarly, we expect that the dissociation points of the excited Efimov states around $a_{*}^{(\mathrm{HL})}/r_{\mathrm{vdw}}=1$-$3$ or $7$-$12$ are more likely to be experimentally observed than those corresponding to the tightly bound states of Table 1. This is analogous to the measurement of $a_{*}$ in systems of three identical bosons, for which the loss-rate peak corresponding to the excited Efimov states appear clearly, whereas those for the ground and first excited states may not appear, or only barely [78, 79].

### v2

@@ II.2 Quantum Defect Theory for non-dipolar system
The QDT parameter $K^{c}$ is formally a matrix acting on multiple channels, but it simplifies as a scalar parameter for the single-channel scattering. Its value is related to the short-range phase, hence the short-range boundary condition. Indeed, at short distance, one can substitute the low-energy form of the wavefunction Eq. (28) because the energy is much smaller than the potential, and imposing $u_{\ell}(r=R_{\mathrm{min}})=0$, one finds
@@
the only difference with Eq. (2) is the presence/absence of the $1/r^{2}$ attraction. At short distance $r\sim R_{\mathrm{min}}\ll r_{\mathrm{vdw}}$, this attraction is much smaller than the van der Waals force so that the two equations are essentially the same. More physically, the short-range phase of the three-body system is dominated by the potential between the heavy atoms. This assumption should be particularly true for a highly mass-imbalanced system where the atomic polarizability of the heavy atoms, hence the van der Waals force and higher-order forces are much stronger than those for the light atom. Therefore, the QDT parameter of the three-body system can be approximated by that of two heavy atoms $K^{c}_{\mathrm{3body}}\simeq K^{c}_{\mathrm{2body}}$, which amounts to adopting the same short-range cutoff $R_{\mathrm{min}}$ value for the two- and three-body systems. While this assumption is plausible for our model, it can be compromised by finite-range effects of the heavy-light atoms’ interaction. It may also be invalidated by the non-adiabatic effects beyond the Born-Oppenheimer approximation, which break down the single-channel description.
@@
Figure 1: Binding energy (in the van der Waals energy unit $1/Mr_{\mathrm{vdw}}^{2}$) of the Efimov states with the unitary heavy-light interaction for (a) bosonic and (b) fermionic heavy atoms interacting via the van der Waals interaction. The mass ratio are (a) $M/m=27.5855...$ and (b) $M/m=27.7520...$, corresponding to 166Er-6Li and 167Er-6Li, respectively. The numerical results for the bosons (blue circles) and fermions (orange squares) with $\ell_{\mathrm{max}}=20$ are shown, together with the analytical curves of Eq. (7) with Eq. (8) (black solid) and Eq. (9) (green dashed). The vertical thin dotted lines (black) denote the thresholds at which the higher angular momentum bound states of $\ell=2$ ($\ell=3$) start to appear for the bosons (fermions), obtained from Eqs. (35) and (8).
@@ III.1 Non-dipolar case $a_{dd}=0$
Figure 2: Binding energy of the Efimov states plotted against (a)(b) the $s$-wave scattering length $a^{(\mathrm{HH})}$ for the bosons, and (c)(d) $p$-wave scattering volume $v^{(\mathrm{HH})}$ for the fermions, obtained by converting $R_{\mathrm{min}}$ in Fig. 1 into $a^{(\mathrm{HH})}$ and $v^{(\mathrm{HH})}$ using Eqs. (8) (13) (14). The left (right) pannels are negative (positive) scattering length/volume sides, which should smoothly connect with each other in the unitary limit $01/a^{(\mathrm{HH})}=$ 0, $01/v_{p}^{(\mathrm{HH})}=$ 0 corresponding to the border between the left and right panels. The solid curves (black) are the analytical results of Eqs. (15) (16), and the dashed curves (green) are the dimer binding energy of the heavy atoms. The vertical dotted lines (black) are the same as those in Fig. 1
First, we show the results without the dipole interaction: in Fig. 1, we show the binding energy of the Efimov states (in the van der Waals energy unit $1/Mr_{\mathrm{vdw}}^{2}$) as the short-range boundary $R_{\mathrm{min}}$ is varied. As $R_{\mathrm{min}}$ gets smaller, the short-range van der Waals attraction increases, leading to the successive appearance of bound states and increase of the binding energy. The numerical data points almost perfectly agree with the analytical curves obtained with the QDT in Eq. (7), particularly at small energy. There are a few points which significantly deviate from the analytical curves. They are ascribed to higher angular momentum states of $\ell=2,4,...$ for the bosons and $\ell=3,5,...$ for the fermions, respectively. The thresholds at which the $\ell=2$ ($\ell=3$) bound states start to appear, denoted by the vertical dotted lines in Fig. 1 (a) (and (b)), agree with most of the stray data points, supporting this interpretation. A few stray points which agree with neither are ascribed to higher angular momentum channels $\ell\geq 4$ ($\ell\geq 5$). This classification of the states has also been confirmed from the shape of the wavefunctions. Because the higher angular momentum states are localized to shorter distances than the $s$-wave ($p$-wave) state owing to the centrifugal barrier, they are more sensitive to $R_{\mathrm{min}}$, resulting in the steeper binding energy curves. The binding energy curves of different partial waves do not render an avoided crossing but simply cross because the angular momentum is a good quantum number in the absence of the dipole interaction.
@@
In Fig. 2, we also show the dimer energy of two heavy atoms (green dashed) obtained by solving Eq. (12). Wherever the Efimov trimers lie above the dimer energy, they are in fact resonant trimers embedded in the dimer-atom continuum [25]. In the Born-Oppenheimer approximation, however, the trimers appear as bound states because the channel describing the dimer-atom continuum is neglected.
@@ III.3 Moderate dipole interaction $a_{dd}\sim r_{\mathrm{vdw}}$
In Fig. 5(g)-(l), we show the binding energies of the Efimov states plotted against $a^{(\mathrm{HH})}_{\mathrm{with-dd}}$. Similar to Fig. 5(a)-(f), the data with different values of $R_{\mathrm{min}}$ mostly collapse into a single curve, demonstrating the universality. One notable difference is that the agreement between the numerical data and the purely van der Waals analytical curves (black) is much better in (g)-(l) than those in (a)-(f): except for the broad avoided crossing at $a_{\mathrm{with-dd}}^{(\mathrm{HH})}/r_{\mathrm{vdw}}\simeq 1$, a significant fraction of the numerical results agree well with the van der Waals curves not only for the weak dipole interaction, but also for the realistic dipole strength (Fig. 5(i)(j)) and stronger dipole interaction ((k)(l)). This suggests that the van der Waals quantum QDT of the bosonic Efimov states in Eq. (15) can capture the universal behavior of the Efimov states in the presence of the dipole interaction, once the effects of the dipole interaction are incorporated into the $s$-wave scattering length between the heavy atoms: $a^{(\mathrm{HH})}=a^{(\mathrm{HH})}_{\mathrm{with-dd}}$. In other words, the major role of the dipole interaction is to change, hence renormalize the value of the $s$-wave scattering length between the heavy atoms, so that the van der Waals universality of the Efimov states in Eq. (15) persists with the renormalized $s$-wave scattering length $a^{(\mathrm{HH})}_{\mathrm{with-dd}}$. This renormalized van der Waals universality for the dipolar bosonic Efimov states we have found here seems non-pertubative; the dipole interaction is so large that it significantly changes both the binding energies (Fig. 5(c)-(f)) and the $s$-wave scattering length (Fig. 6(b)(c))) from the purely van der Waals system. Some horizontal data points in Fig. 5(g)(l) which are not present in (a)-(f) are artefacts of the higher-partial-wave induced resonances: the two-body problem between the heavy atoms in Eq. (17) not only shows the $s$-wave dominated resonances, but also shows the resonant enhancement of $a^{(\mathrm{HH})}$ induced by higher-partial-wave channels [59, 61], whose approximate positions are indicated by the vertical black dotted lines in Fig. 6. The narrow nature of these higher-partial-wave induced resonances leads to an abrupt change of $a^{(\mathrm{HH})}_{\mathrm{with-dd}}$ as $R_{\mathrm{min}}$ is changed, resulting in flat energy spectra in Fig. 5(g)-(l). Equation (15) cannot explain these flat data points. We note however that these flat energy spectra are absent in Fig. 5(a)-(b), and they should therefore be interpreted as artefacts originaing from the two-body effect of $a^{(\mathrm{HH})}$ rather than a genuine three-body phenomenon. Physically, when we attempt to evaluate the binding energies of the Efimov states from an experimentally obtained value of $a^{(\mathrm{HH})}$ (see Sec. IV), we can simply use the van der Waals curves of Eq. (15) by neglecting the stray flat spectra, except for the case where the two Er atoms are coincidentally in the proximity of the $s$-wave resonance induced by a $d$-wave or higher-angular-momentum molecular state. In the avoided crossing region, the variance of the spectrum is larger than in the other regions (see Fig. 5(l)). This is due to the finite values of $R_{\mathrm{min}}$, as it can be checked that reducing it further does lead to a converged universal behavior. For the finite values of $R_{\mathrm{min}}$ used in the figures (which is varied between 0.25 and 0.4 $r_{\mathrm{vdw}}$), it is still justified to call this region’s behavior universal despite the noticeable energy variance because it remains much smaller than the Efimov period, in sheer contrast with the strength of the van der Waals interaction at short distance which varies by a factor 16. Yet, a $d$-wave dimer of the two heavy atoms appearing around $a_{\mathrm{with-dd}}^{(\mathrm{HH})}/r_{\mathrm{vdw}}\simeq 1$ (see Fig. 11 in Appendix. D where the dimer energies are presented) seems to significantly affect two-body and three-body behaviors in this region, making the simple renormalized van der Waals universality of Eq. (15) inadequate.
@@
In Fig. 8, we show the binding energies of the fermionic Efimov states plotted against $K_{c,F}^{(3D)}$ as obtained above. The numerical data of different short-range $R_{\mathrm{min}}$ values collapse into a single curve, demonstrating the universality of the fermionic Efimov states. The universality holds particularly well for a realistic $a_{dd}$ value of 167Er-167Er (Fig. 8(a)(b)), where only a few points of higher angular momentum character deviate from the universal curve. In contrast to the bosonic case, we cannot find an analytical expression which could explain the universal energy spectra owing to the impossibility of using $v_{p}^{(\mathrm{HH})}$ hence Eq. (16). The avoided crossing at $K_{c,F}^{(3D)}\simeq-1$ originates from the $\ell=3$ channel which is present without the dipole interaction at around $r_{\mathrm{vdw}}/v_{p}^{(\mathrm{HH})}\simeq-0.3$ (see Fig. 2(c) and Fig. 12 in Appendix D). It shifts toward a larger negative $K_{c,F}^{(3D)}$ side with a decreased slope as the dipole strength is increased from a realistic value of 167Er-167Er (Fig. 8(a)(b)) toward larger values (Fig. 8(c)-(f)). The foothill of the avoided crossing shows a larger variance for stronger dipole interactions (Fig. 8(c)-(f)). This is likely be a precursor of the breakdown of our assumption that the single-channel short-range phase in $\ell=1$ universally characterizes the three-body system (see Sec. III.4), especially for a very strong dipole interaction and large $K_{c,F}^{(3D)}$ (Fig. 8(c)(e)(f)).
@@ III.4 Strong dipole interaction $a_{dd}\gg r_{\mathrm{vdw}}$
The data points with a small slope in Fig. 9, including the weakly bound states, tend to collapse into a single curve for both the bosonic and fermionic systems. Although the spreading of the data points is larger than those in Figs. 5 and 8, it remains much smaller than the Efimov period, even with an almost 16 times change in the short-range part of the van der Waals interaction. We have also confirmed that the spreading gets smaller as $R_{\mathrm{min}}$ is decreased, in a similar manner to Fig. 5. We can thus conclude that $K_{c}^{(1D)}$ universally characterizes the short-range phase, hence the Efimov states with the strong dipole interactions. On the other hand, the data points with steeper slopes do not seem to show the universality with respect to $K_{c}^{(1D)}$. Since those states have significant contributions from large angular momentum states $\ell\gg 1$ and the angular insensitive nature of their short-range phase can break down, it is not unexpected to find that their features cannot be universally captured by the one-dimensional scattering model in Eqs. (20) (21) where the angular momentum is absent (see also dimers in Fig. 13 in Appendix D with similar behaviors). Nevertheless, the physical mechanics behind the universal and non-universal behaviors in Sec. III.3 and Fig. 9 and how they change in-between them have yet to be clarified.
@@ IV Three-body parameters of Er-Er-Li Efimov states
166Er-6Li 75.5 65.5 68 0.495 -10.1 0.269
0.107 -46.7 1.25
2.29×$\times 10^{-2}$217 5.81
4.94×$\times 10^{-3}$1.01×$\times 10^{3}$ 27.0
168Er-6Li 75.8 66.3 137 0.352 -14.2 0.381
7.65×$\times 10^{-2}$65.3 1.75
1.66×$\times 10^{-2}$300 8.05
3.62×$\times 10^{-3}$1.38×$\times 10^{3}$ 37.0
170Er-6Li 76.0 67 221 0.298 -16.8 0.452
6.55×$\times 10^{-2}$76.6 2.06
1.44×$\times 10^{-2}$349 9.37
3.16×$\times 10^{-3}$1.59×$\times 10^{3}$ 42.7
While there are spectroscopic methods to directly associate the Efimov trimers and observe their binding energies [74, 75, 76, 77], it is challenging to perform such experiments, especially in the unitary limit. It is more straightforward to observe the three-body loss rate of the atoms for variable $s$-wave scattering lengths $a^{(\mathrm{HL})}$, and search for the negative scattering length $a=a_{-}^{(\mathrm{HL})}$ at which the loss rate peak appears [11, 12, 13, 14, 16, 19]. Alternatively, one can prepare a mixture of heavy-light Feshbach molecules and heavy atoms, and observe the dimer-atom loss rate to find the positive scattering length $a=a_{*}^{(\mathrm{HL})}$ at which its peak appears [17]. The values $a_{-}^{(\mathrm{HL})}$ and $a_{*}^{(\mathrm{HL})}$ correspond to the points at which the Efimov states dissociate into three atoms, or a dimer and an atom, respectively. Notably, any of $a_{-}^{(\mathrm{HL})}$, $a_{*}^{(\mathrm{HL})}$, or $\kappa_{*}$ can be regarded as the universal scale characterizing the Efimov states because they can be related with each other by the universal Efimov theory [1, 2, 3, 4, 5, 6]. Using the zero-range Efimov theory, we have estimated $a_{-}^{(\mathrm{HL})}$, $a_{*}^{(\mathrm{HL})}$ from the $\kappa_{*}$ values. The results are presented in the last two columns of Table 1. The values of $a_{-}^{(\mathrm{HL})}$ corresponding to the most tightly bound Efimov states are likely to be either around $a_{-}^{(\mathrm{HL})}/r_{\mathrm{vdw}}\approx$ $-10$-$17$, or adjacent ones with $\approx-2$-$4$ (not tabulated). With the low-energy formula Eq. (15) and the zero-range theory, we cannot conclude which of them are the ground Efimov state. We also note that they can be significantly affected by finite-range effects owing to their relatively large binding energy, while results for the more weakly bound states in Table 1 should be more quantitatively reliable. Similarly, we expect that the dissociation points of the excited Efimov states around $a_{*}^{(\mathrm{HL})}/r_{\mathrm{vdw}}$$\approx 5$-$9$ are more likely to be experimentally observed than those of $\approx 1-2$ or $\approx 0.2-0.4$ corresponding to the tightly bound states. This is analogous to the measurement of $a_{*}$ in systems of three identical bosons, for which the loss-rate peak corresponding to the excited Efimov states appear clearly, whereas those for the ground and first excited states may not appear, or only barely [78, 79].